OpenStudy (anonymous):

Solve for x by using the quadratic formula 2x2+9x-4=0

OpenStudy (abb0t):

Quadratic Formula: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (abb0t):

Where: \[ax^2 + bx +c\]

OpenStudy (anonymous):

ok i got \[-9\pm \sqrt{113}\div4\] but what do i do futher from there?

OpenStudy (abb0t):

You should leave it as \[\frac{ -9 \pm \sqrt{113} }{ 4 }\] that is the exact answer, but if you want to get a numerical value, you can add -9+√113 and divide by 4 AND -9 - √113 and divide by 4. You will get two values for x.

OpenStudy (anonymous):

Thank you